Fermi National Laboratory
 

Operation Summary for February 24, 2010 - February 26, 2010

 

Wednesday Feb. 24

The day shift began with Operations monitoring the Tevatron (TeV) store 7631 (133.59E30) and the Recycler stash (105.2E10), with the Antiproton Source stacking (7.6E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Linac RF personnel investigated station LRF5; they replaced a series of failed SCRs. Booster RF personnel accessed the tunnel and replaced various RF power amplifiers. Operations resumed sending beam to Pbar, NuMI, and MiniBooNE at 11:25 AM.

During the evening shift, Operations terminated store 7631 at 10:50 PM and put the TeV into HEP shot setup.

Thursday Feb. 25

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (13.1E10), with Pbar stacking (43.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established store 7632 at 12:25 AM with an Average Initial Luminosity (AIL) of 288.92E30. Operations reported problems with BRF14. Operations established beam to MTest at 5:05 AM. At 6 AM, the Meson Lambertson power supply (MLAM2) tripped off due to a magnet temperature indication; Operators accessed enclosure C to verify that the LCW valves to the Lambertson were open; they were. Operations reestablished beam to MTest at 7:30 AM.

During the day shift, Operations terminated store 7632 at 2:05 PM and then prepared CDF and D0 for accesses into their collision halls. Operators rebooted the MI low-level RF (LLRF) system to load updated code. Operators prepared the TeV A-E enclosures for access to investigate a magnet heater issue. They came out of the TeV access at the end of the shift.

During the evening shift, Operations began the TeV turn on at 4:30 PM. The new MI LLRF code had caused problems with transfers of antiprotons to the Recycler so the previous version of code was reloaded. The Time Line Generator (TLG) locked and inhibited the loading of new timelines. Operations contacted Controls personnel and the TLG crate was rebooted. Operations put the TeV into shot setup at 8:12 PM and established store 7634 at 9:25 PM with an AIL of 317.43E30.

Friday Feb. 26

The midnight shift began with Operations monitoring store 7634 (173.84E30) and the stash (106.6E10), with Pbar stacking (23.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, a Linac Klystron RF station (KRF1) tripped off at 12:05 AM due to a PFN door open interlock indication. Operators called in Linac personnel to investigate and they had the beam restored by 4:15 AM. BRF1 tripped off and required a reduction in Booster intensities for all beam cycles.

The Plans for Friday and the Weekend

The plans for today are to allow Booster to resolve their RF problems. The Run Coordinator will also allow Linac, MI, and Pbar personnel to conduct necessary maintenance work that doesn’t interfere with the stack, stash, or store. The plans for the rest of the day and weekend will consist of stacking, stashing, and storing.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/26/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory